### v2

# Exceptional topology in Non-Hermitian Twisted Bilayer Graphene

Abstract
Twisted bilayer graphene has extraordinary electronic properties at the magic angle along with an isolated flat band at magic angle. However, the non-Hermitian phenomena in twisted bilayer graphene remain unexplored. In this work, we study a non-Hermitian TBG formed by one-layer graphene twisted relative to another layer with gain and loss. Using a non-Hermitian generalization of Bistritzer-MacDonald model, we find Dirac cones centered at only $K_{M}$ ($K^{\prime}_{M}$) corner of the moiré Brillouin zone at $K^{\prime}$ ($K$) valley deformed in the presence of non-Hermiticity. This is different from single layer graphene with gain and loss, where rings of exceptional points appear in both $K$ and $K^{\prime}$ corners of the Brillouin zone. The coincident of exceptional rings at $\Gamma_{M}$ point characterizes an “exceptional magic angle”, at which the system hosts flat bands with zero energy and finite lifetime. More interestingly, we find that the topological charge in the moiré Brillouin zone is conserved during the expansion and fusion of the exceptional ring, which is absent in two-dimensional systems constraining by Nielsen-Ninomiya theorem. These findings can be demonstrated in realistic cold atom and metamaterial systems and will stimulate further study on non-Hermitian phenomena in twistronic.
I introduction
Figure 1: (color online)Schematics of twisted bilayer graphene with onsite gain and loss. (a) Real-space geometrical picture: The gray (orange) honeycomb lattice marks the top (bottom) layer. The red (blue) dots represent the amplifying (lossy) sites in the bottom layer. (b) Momentum-space geometrical picture: The gray (brown) lattice marks the Brillouin zone of the top (bottom) layer graphene. Their expansion around the $K$ ($K^{\prime}$) valley forms moiré Brillouin zone in green (blue). (c) Moiré Brillouin zone physics at $K^{\prime}$ valley: Under the effect of onsite gain and loss, Dirac points at $K_{M}$ corners (related to $K_{b}$ points of bottom layer graphene) morph into exceptional rings. At an “exceptional magic angle”, the exceptional rings touch at $\Gamma_{K}$ point. (d) Moiré Brillouin zone physics at $K$ valley: exceptional rings from Dirac points at $K^{\prime}_{M}$.
Graphene owes extraordinary electronic and optical properties in two-dimensional (2D) systems  [1, 2]. These properties relate to its lattice structure, which is directly interconnected with its topological property. In single-layer graphene, the presence of unusual linear band dispersion termed “Dirac cone” at $K$ and $K^{\prime}$ points of the Brillouin zone is derived from its honeycomb lattice geometry. The two Dirac points are related by time-reversal symmetry and thus have opposite Berry phases, which has been confirmed in experiments [3, 4].
When two graphene layers are shifted with each other in rational competing periodicities upon shearing or twisting, they form a moiré pattern [5, 6]. One prominent example is a graphene bilayer with a relatively small angle rotation between the layers [7]. The recent studies on twisted bilayer graphene (TBG) focus on the extraordinary electronic properties at the magic angle along with an isolated flat band [8, 9, 10, 11, 12]. It is related to a variety of exotic phases including quantum anomalous Hall [13, 14, 15, 16, 17], fractional Chern insulators [18, 19, 20], ferromagnetic states [14, 21, 22], nematicity [23], unconventional superconductivity [24, 25, 26, 27, 28, 29]. On the other hand, twisted bilayer graphene has extraordinary topological properties. Since the two Dirac points in moiré Brillouin zone (mBZ) in one valley emanate from different layers, their Berry phases can be identical due to the symmetry interlayer hopping term, which is generally not allowed in two-dimensional periodic systems following the Nielsen-Ninomiya theorem [30, 31]. This is in contrast to single-layer graphene, in which the Berry phases of Dirac cones in different valleys has different topological charges.
The topological properties of two-dimensional systems can be generalized to non-Hermitian systems by introducing on-site gain and loss [32, 33, 34]. The presence of non-Hermiticity can transform Hermitian degenerate points, such as a Dirac-like point or Weyl point, into a ring of exceptional points [35, 36]. At an exceptional point, not only the real and imaginary parts of the eigenvalues degenerate, but also the eigenvectors coalesce. Many phenomena are unique to non-Hermitian systems, including non-Hermitian skin effect [37] and non-Hermitian topological classifications [38, 39]. Despite topological robustness has been shown in one-dimensional moiré lattice under stain, the non-Hermitian effect of twisted moiré system with magic angle has been overlooked [40, 41].
Usually, non-Hermiticity is detrimental to the topology in single-layer graphene. When the exceptional contours with opposite topological charges merge, the topological charge can be dissipated [36]. In single-layer graphene, the annihilation of exceptional points from different valleys has been experimentally observed in a two-dimensional photonic system [42]. Since the moire BZ valley of TBG has different topological properties from single-layer graphene, it is natural to ask how the exceptional geometry appears and develops in a non-Hermitian TBG system.
In this paper, we uncover a new type of twisted band engineering in TBG, with non-Hermitian perturbation on one of the layers. As shown in Fig. 1(a), balanced gain and loss are put on the bottom layers (in red and blue, respectively). The superposing of two graphene layers with a twist angle between them creates moiré patterns. In the mBZ shown in Fig. 1(b), the low-energy band structure consists of Dirac cones from the rotated layers located at the $K_{M}$ and $K^{\prime}_{M}$ corners. When the non-Hermitian perturbation is turned on the bottom layer, Dirac points at $K_{M}$ ($K^{\prime}_{M}$) corner in the mBZ at $K^{\prime}$ ($K$) valley morph into exceptional rings and the corresponding Dirac cones become exceptional cones, as shown in Fig. 1(c) and (d) respectively. We find that this non-Hermitian system hosts special angles for which multiple exceptional cones coincide at $\Gamma_{M}$ points in an mBZ. We call this phenomenon “exceptional magic”. And importantly, our results demonstrated that the total Berry charge is conserved in the mBZ even after the merging of exceptional rings, this is a large difference from non-Hermitian single-layer graphene.
The paper is organized as follows. Section II introduces the non-Hermitian TBG model. Both the real and imaginary parts of the low-energy moiré band spectrum will be explored in Sec. III. In Sec. IV, we will use a non-Hermitian three triod model to discuss exceptional physics. In Sec. V, we calculate Beery curvature and the corresponding Chern number. Finally, we conclude our results in Sec. VI.
II The non-Hermitian bilayer model
We consider balanced gain and loss on the bottom layer of the bilayer system. We can write down a non-Hermitian generalization of the single-valley model for twisted bilayer graphene. The low-energy Hamiltonian for the mBZ at $K$ or $K^{\prime}$ valley is:
$H=\begin{pmatrix}h_{b}(\bf{k})&\bf{T}(\bf{r})\\ \bf{T}(\bf{r})^{\dagger}&h_{t}(\bf{k})\end{pmatrix}$ (1)
where the top layer Hamiltonian is
$h_{t}=-iv_{F}\bm{\sigma}_{\theta/2}\bm{\nabla}$ (2)
and the bottom layer Hamiltonian is non-Hermitian
$h_{b}=-iv_{F}\bm{\sigma}_{-\theta/2}\bm{\nabla}+i\lambda_{V}\sigma_{z}.$ (3)
Here, $i\lambda_{V}\sigma_{z}$ is a balanced gain and loss imposing on the bottom layer, which induces non-Hermiticity. And $\bm{\sigma}_{\theta/2}$ corresponds to the rotated Pauli matrices $e^{-i\theta\sigma_{z}/4}(\pm\sigma_{x},\sigma_{y})e^{i\theta\sigma_{z}/4}$ with $\pm$ for $K$ and $K^{\prime}$ valley respectively. $\textbf{T}(\textbf{r})=\sum^{3}_{j=1}T_{j}e^{-ik_{\theta}\textbf{q}_{j}\cdot\textbf{r}}$ is the sublattice-dependent moiré potential that couples the two layers. The interlayer coupling is in the form
$T_{j+1}=w_{0}\sigma+w_{1}(\sigma_{x}\cos j\phi+\sigma_{y}\sin j\phi)$ (4)
with $\phi=2\pi/3$ and $w_{0}(w_{1})$ the interlayer coupling strength in the AA (AB) region. The unit vectors $\textbf{q}_{1}=(0,-1)$, $\textbf{q}_{2,3}=(\pm\sqrt{3}/2,1/2)$. $k_{\theta}=2k_{D}\sin(\theta/2)$ is the moiré modulation vector and $k_{D}=4\pi/(3a_{0})$ is the magnitude of the Dirac wave vector, where $a_{0}$ is the lattice constant of monolayer graphene. The hopping strength $w$ encodes the inter-layer coupling in the AB region.
The generators of the magnetic space group $P6^{\prime}2^{\prime}2$ for the Hermitian twisted bilayer graphene include a $C_{3z}=e^{i\frac{2\pi}{3}\sigma_{z}}$ rotation symmetry, a $C_{2x}=\tau_{x}\sigma_{x}$ rotation symmetry, and a $C_{2z}T=\sigma_{x}\mathcal{K}$ symmetry [43]. The single-valley Hamiltonian only has the $C_{3z}$ symmetry, since $C_{2z}$ and $T$ symmetries play the role of mapping one valley to the other. We can check that the non-Hermitian model preserves the $C_{3z}$ rotation symmetry.
If there is no AA stacking ($w_{0}=0$), the Hamiltonian preserves chiral symmetry $\mathcal{C}\mathcal{H}(\textbf{r})\mathcal{C}^{\dagger}=-\mathcal{H}(\textbf{r})$ under chiral symmetry operator $\mathcal{C}=\sigma_{z}$ in the absence of non-Hermiticity [44, 45, 46, 43].
In the presence of the balanced gain and loss $i\sigma_{z}$ term, the chiral symmetry is broken but non-Hermitian chiral symmetry is preserved In the following numerical calculation, we take $\hbar v_{F}/a=2380meV$, $w_{1}$ = 110meV and $2v_{0}k_{D}=19.81eV$. They correspond to the first magic angle $\alpha\approx 0.586$.
III The appearance and development of exceptional rings
Figure 2: (Color online) (a) The real part and (b) the imaginary part of the band spectra in the mBZ for $\theta=1,u_{0}=0,u_{1}=0.005,\lambda_{V}=0.05$.
Figure 3: (color online) (a) The real part and (b) the imaginary part of the four lowest bands in the mBZ for $u_{0}=0,u_{1}=0.005,\lambda_{V}=0.072$. The real parts of the two lowest bands are flat.
Let us investigate the band structure for twist angle $\theta=1$. In the absence of non-Hermiticity, this is not a magic angle with a flat band. If the interlayer coupling effect is absent ($w_{0}=w_{1}=0$), the dispersion in the mBZ is equivalent to the folded dispersion in the Brillouin zone of graphene. Since the physics in $K$ valley and $K^{\prime}$ valley are related, we only discuss the $K$ valley in the following. The Dirac point at the $K$ corner in the top layer is mapped to $K_{M}$ point in the mBZ, while that at the $K$ corner of the bottom layer is mapped to $K^{\prime}_{M}$ point in the mBZ.
In the presence of non-Hermitian perturbation $\lambda_{V}$, the energy dispersion becomes complex. In Fig. 2(a), we can see that non-Hermitian perturbation deforms the Dirac cones at $K^{\prime}_{M}$ corners. More specifically, the degeneracy point at the $K^{\prime}_{M}$ points in the real part of the spectrum morphs into a ring of degeneracy points. This is a stark difference from single-layer graphene with onsite gain and loss. The latter has deformed Dirac cones in both $K$ and $K^{\prime}$ corners in the Brillouin zone. This difference is due to the fact that the Dirac cones at $K_{M}$ and $K_{M}^{\prime}$ point in non-Hermitian TBG are folded from different graphene layers. The non-Hermitian $\lambda_{V}$ can only deform the Dirac cone on the bottom layer.
In particular, the degeneracy of the two lowest energy states leads to the appearance of singularity points in momentum space. At the same time, a ring of degeneracy point appears at the same momentum position in the imaginary part of the spectrum, as shown in Fig. 2(b). We found that not only the eigenvalues coalesce but also their corresponding eigenvectors coalesce at this point. This confirms that these singularity points are exceptional points in momentum space. Thus, the Dirac cone is turned into a ring of exceptional points, this is similar to that in non-Hermitian semimetals [35].
At a critical non-Hermitian strength $\lambda_{Vc}$, the exceptional point approach $\Gamma_{M}$ point, where the real part whole lowest band becomes flat, as shown in Fig. 3. This characterizes an “exceptional magic” angle. According to the geometry relationship, $\lambda_{Vc}$ is determined by the lattice size of mBZ $k_{\theta}$ and Fermi velocity $v_{F}$.
IV Exceptional cone
To understand the nonhermitian physics, we use the simplified tripodmode model [8], which truncates Eq. 1 at the first honeycomb shell. The Hamiltonian is written as
$H_{tri}=\begin{pmatrix}h_{b}(\bf{k})&T_{1}&T_{2}&T_{3}\\ T_{1}&h_{t1}(\bf{k}-\bf{q}_{1})&0&0\\ T_{2}&0&h_{t2}(\bf{k}-\bf{q}_{2})&0\\ T_{3}&0&0&h_{t3}(\bf{k}-\bf{q}_{3})\end{pmatrix}$ (5)
where $h_{tj}$ for $j=1,2,3$ are on the top layer and $h_{b}$ on the bottom layer. Since $h_{b}$ is non-Hermitian, the Hamiltonian $H_{tri}$ has biorthogonal eigenvectors. In this system, the left and right eigenvectors are four two-component spinors
$\Psi_{\alpha}^{T}(\textbf{k})=(\psi_{0}(\textbf{k}),\psi_{1}(\textbf{k}),\psi_{2}(\textbf{k}),\psi_{3}(\textbf{k}))_{\alpha}^{T},\quad\alpha=L,R.$ (6)
Using perturbation theory one can derived the effective Hamiltonian fin the space of $\psi_{0}$ The Hamiltonian $H_{tri}$ can be divide into a momentum-independent part $\mathcal{H}^{(0)}$ and a momentum-dependent part $\mathcal{H}^{(1)}_{k}$. The $\psi_{jR}$ can be expressed by
$\psi_{jR}=-h_{j}^{-1}T_{j}\psi_{0R}.$ (7)
Since $h_{j}$ is independent of non-Hermicity, we can expect $T_{j}h_{j}^{-1}T_{j}=0$, which is similar to the Hermitian model. Thus, the $\psi_{0}$ spinor in the zero-energy eigenstates satisfies
$h_{0}\psi_{0R}=0.$ (8)
This form seems to be the same as that of the Hermitian TBG. However, $h_{0}$ is in a non-Hermitian form, the $\psi_{0R}$ is different from $\psi_{0L}$. We can check that the biorthorgonal wave functions are normalized as
$\langle\Psi_{0L}|\Psi_{0R}\rangle=1+3(\alpha_{0}^{2}+\alpha_{1}^{2}).$ (9)
The effective Hamiltonian matrix to leading order in $k$ is
$\langle\Psi^{(i)}|\mathcal{H}^{(1)}_{\textbf{k}}|\Psi^{(j)}\rangle=\psi_{0}^{(i){\dagger}}(-v_{F}^{*}\rm{\sigma}\cdot\mathbf{k}+i\lambda_{V}^{*}\sigma_{z})\psi^{(j)}_{0}$ (10)
with the renormalized velocity being
$\frac{v_{F}^{*}}{v_{F}}=\frac{1-3\alpha_{1}^{2}}{1+3(\alpha_{0}^{2}+\alpha_{1}^{2})}$ (11)
and the renormalized non-Hermitian strength being
$\frac{\lambda_{V}^{*}}{\lambda_{V}}=\frac{1-3(\alpha_{0}^{2}-\alpha_{1}^{2})}{1+3(\alpha_{0}^{2}+\alpha_{1}^{2})}.$ (12)
We can see that aside from a renormalized velocity and non-Hermitian strength, the form of Hamiltonian is identical to the continuum model Hamiltonian of non-Hermitian single-layer graphene. The form of the renormalized velocity is the same as that of the Hermitian TBG. It decreases with increasing interlayer coupling in AB region $\alpha_{1}$. To see the effect of interlayer coupling on non-Hermitian strength, we should discuss two cases for $\alpha_{0}=0$ and $\alpha_{0}\neq 0$. When $\alpha_{0}=0$, $\lambda_{V}^{*}$ increases with increasing interlayer coupling in AB region $\alpha_{1}$. Thus, the ratio $\lambda_{V}^{*}/v_{F}^{*}$ becomes larger as $\alpha_{1}$ increases. When $\alpha_{0}\neq 0$, $\lambda_{V}^{*}$ increases with increasing interlayer coupling in AA region $\alpha_{0}$ and the non-Hermitian effect is more complicated. We can expect that the exceptional magic angle $\alpha_{1c}$ is not only determined by $\lambda_{V}$ but also $\alpha_{0}$.
V Berry Curvature and Chern number
So far, we have shown the existence of exceptional geometries in the non-Hermitian TBG. In this section, we will explore their topological properties numerically.
The topological charge or Chern number of an exceptional contour can be obtained by integrating the Berry connection $A(\mathbf{k})$ along a closed surface containing the exceptional contour
[…]
where the Berry connection is defined by the eigenvectors as
$\mathbf{A}^{\beta,\beta^{\prime}}=i\langle\psi^{\beta}(\mathbf{k})|\nabla_{\mathbf{k}}|\psi^{\beta^{\prime}}(\mathbf{k})\rangle.$ (14)
Here, $|\psi^{\beta^{\prime}}(\mathbf{k})\rangle$ and $\langle\psi^{\beta}(\mathbf{k})|$ can be the left or right eigenvectors. Since previous study has proven that the total topological charge is the same for four possible choices of eigenvectors [47], we will choose $\beta=L$ and $\beta^{\prime}=R$ in the following calculation.
Figure 4: (color online) Berry curvature for the $\sigma_{z}=1$ band in the moiré Brllouin zone for $u_{0}=0$, $u_{1}=0.005$, $\theta=1$ and $\lambda_{V}=0.08>\lambda_{Vc}$. A small in-plane electric field is imposed to open the gap.
In numerical calculation, the Berry connection is presented by the local Berry curvature $\Omega(\mathbf{k})=\nabla_{\mathbf{k}}\times\mathbf{A}(\mathbf{k})$, which can be calculated as
$\Omega(\mathbf{k})^{\beta,\beta^{\prime}}=\lim_{q\rightarrow 0}\frac{1}{4q^{2}}\langle\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}-q\hat{y}}|\psi^{\beta}_{\mathbf{k}-q\hat{x}+q\hat{y}}\rangle\langle\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}+q\hat{y}}|\psi^{\beta}_{\mathbf{k}+q\hat{x}+q\hat{y}}\rangle\langle\psi^{\beta}_{\mathbf{k}+q\hat{x}+q\hat{y}}|\psi^{\beta^{\prime}}_{\mathbf{k}+q\hat{x}-q\hat{y}}\rangle\langle\psi^{\beta}_{\mathbf{k}+q\hat{x}-q\hat{y}}|\psi^{\beta^{\prime}}_{\mathbf{k}-q\hat{x}-q\hat{y}}\rangle$ (15)
with $q$ being half of the lattice constant.
Figure 4 shows the Berry curvature in mBZ. We can see that Berry curvatures have sharp peaks around $\Gamma_{M}-K_{M}$ high symmetric lines, which are close to the boundary of exceptional rings. This can be well understood since Berry curvature is the phase of an overlap tracing a small loop in mBZ. On the boundary of exceptional rings, the wave functions vary rapidly and will give a large Berry curvature.
Upon summing the Berry curvature on the mBZ, we can obtain the total Berry charge or Chern number
$\mathcal{C}=\sum_{mBZ}\Omega^{LR}(\textbf{k}).$ (16)
The non-Hermitian perturbation transforms a Dirac point into an exceptional ring. In the mBZ, there are three-thirds of exceptional rings from Dirac points in different moiré corners. For small $\lambda_{V}$, we can find that the total Berry charge in one mBZ is 1. This can be easy to understand from two facts. One is that each of the three Dirac points corresponds to Berry curvature monopoles of charge 1 in the absence of non-Hermiticity. The other is that the topological charge is preserved on the exceptional contour that forms from the original Dirac point. Thus, we can expect that each of the three exceptional contours contributes one-thirds of the Berry charge to one mBZ.
Most interestingly, even when $\lambda_{V}$ exceeds $\lambda_{Vc}$, the total Berry charge remains one. This is opposite to the merging of exceptional contours with opposite topological charge [36], in which the topological charge is dissipated and gives a single, uncharged exceptional contour. This indicates a breakdown of Nielsen-Ninomiya theorem and distinguishes the valley topological features of non-Hermitian TBG from a single-layer graphene.
At much larger $\lambda_{V}$ (e.g., $\lambda_{V}>0.85$ for $\theta=1$), the Berry charge can be a quantized value larger than 1 due to the band crossings with the remote bands.
VI Discussion and conclusion
In this paper, we have studied a non-Hermitian generalization of twisted bilayer graphene system with balanced gain and loss on one of the layers. We found that exceptional rings centered at $K_{M}$ ($K^{\prime}_{M}$) corner in the moiré Brillouin zone in the $K^{\prime}$ ($K$) valley. This is due to the folding of the Brillouin zone, and it is absent in the single-layer graphene. The exceptional rings coincide and form bands whose real parts are flat at a specific non-Hermitian strength, which is denoted as “exceptional Dirac magic”. This flat band is characterized with a robust Chern number $\mathcal{C}=1$ in moiré Brillouin zone. This is different from single-layer graphene in which Dirac/Weyl exceptional rings carrying opposite topological charges are coupled and annihilated.
We have introduced a new design concept for twisted moiré systems, in which a certain “exceptional Dirac magic” angle is achieved. Non-Hermitian TBG can be realized in cold atom and metamaterial system. On one hand, due to the great advantage of manipulating atom-atom interaction and light-matter coupling, ultracold atom experiments provide an unprecedented opportunity. Remarkably, atomic Bose-Einstein condensate has been recently realized in twisted bilayer optical lattices  [48]. Also, there are proposals on the photonic analog [49, 50, 51, 52, 53] and phononic analog [54, 55] of twisted bilayer graphene. In particular, there is demonstration of optical bilayer photonic crystal devices in the microwave range [56] and in the optical frequency range [57]On the other hand, on-site dissipations are well-controlled non-Hermitian terms that can be realized experimentally in many systems [32, 33, 34]. In particular, the Weyl exceptional ring has been realized in photonic experiment [58, 59].
The highlight of this work is the discovery of non-dissipation exceptional topology. The interplay of topology, non-Hermitian, and moiré physics can open the door to new phenomena. Especially, the reduced kinetic energy of the flat bands in twisted bilayer graphene makes it a correlated system. Non-Hermitian strongly interacting Dirac system has recently been calculated [60]. The non-Hermitian physics in quantum many-body systems, such as strongly correlated phases in moié system, will be a future focus.
Note added-. Upon completion of our manuscript, we became aware of a recent paper on non-Hermitian twisted bilayer graphene, in which non-Hermiticity is on the hopping amplitude and topological aspects are undiscussed [61].
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No. 12104099 and No. 12274095) and Guangzhou Science and Technology Program (Grant No. 2024A04J0272)